OpenStudy (anonymous):

show all work. Last fall, a gardener planted 95 iris bulbs. She found that only 80 of the bulbs bloomed in the spring. Find the empirical probability that an iris bulb of this type will bloom. Give answer as a fraction in lowest terms. How many of the bulbs should she plant next fall if she would like at least 91 to bloom? please show work!

OpenStudy (anonymous):

\[\frac{80}{95}=\frac{16}{19}\] in lowest terms

OpenStudy (anonymous):

for the second part solve \[\frac{16}{19}=\frac{91}{x}\] for x \[x=\frac{19\times 91}{16}\] or about 108

OpenStudy (anonymous):

since it says "at least" you might want to say 109

OpenStudy (anonymous):

please help me!!

OpenStudy (anonymous):

what step?

OpenStudy (anonymous):

im not understanding what u wrote

OpenStudy (anonymous):

can't you read it or don't you know where i got the numbers from?

OpenStudy (anonymous):

if it looks funny refresh your browser because i wrote it in latex. you should be able to see the fractions clearly